A space craft flying in a straight course with a velocity 75km/s fires it's rocket motors for 6.0 seconds . At the end of this time it's speed is 120 km/s in the same direction. Find(i) the space craft's average accleration while the motors are firing, (ii) the distance travelled by the space craft in the first 10 seconds after the rocket motors having been in action for only 6.0s .
Plzzzz solve. .......
It would be a great help indeed !!!!!!!
Thank you

Answer by Alan3354(69443)

A space craft flying in a straight course with a velocity 75km/s fires it's [sic] rocket motors for 6.0 seconds. At the end of this time it's [sic] speed is 120 km/s in the same direction. Find(i) the space craft's average accleration while the motors are firing,
Change in speed = 120 - 75 = 45 km/sec
Accel = 45/6 = 7.5 km/sec/sec
======================
(ii) the distance travelled by the space craft in the first 10 seconds after the rocket motors having been in action for only 6.0s
---
If you mean after the 6 second firing:
d = r*t = 120*10
d = 1200 kms
---
If you mean 10 starting at the 6 second firing:
The average speed for the 6 seconds = (75 + 120)/2 = 97.5 km/sec
d = 97.5*6 + 4*120
d = 1065 kms
-----------------
PS it's = it is